# On the Evaluation of Generative Models in Distributed Learning Tasks

### v3

@@ Abstract
The evaluation of deep generative models including generative adversarial networks (GANs) and diffusion models has been extensively studied in the literature. While the existing evaluation methods mainly target a centralized learning problem with training data stored by a single client, many applications of generative models concern distributed learning settings, e.g. the federated learning scenario, where training data are collected by and distributed among several clients. In this paper, we study the evaluation of generative models in distributed learning tasks with heterogeneous data distributions. First, we focus on the Fréchet inception distance (FID) and consider the following FID-based aggregate scores over the clients: 1) FID-avg as the mean of clients’ individual FID scores, 2) FID-all as the FID distance of the trained model to the collective dataset containing all clients’ data. We prove that the model rankings according to the FID-all and FID-avg scores could be inconsistent, which can lead to different optimal generative models according to the two aggregate scores. Next, we consider the kernel inception distance (KID) and similarly define the KID-avg and KID-all aggregations. Unlike the FID case, we prove that KID-all and KID-avg result in the same rankings of generative models. We perform several numerical experiments on standard image datasets and training schemes to support our theoretical findings on the evaluation of generative models in distributed learning problems.
Generative Model, Evaluation, Distributed Learning
@@ 1 Introduction
Deep generative models including diffusion models (Sohl-Dickstein et al., 2015) and generative adversarial networks (GANs) (Goodfellow et al., 2014) have attained impressive results over a wide array of machine learning tasks (Karras et al., 2019; Ho et al., 2020; Ramesh et al., 2022). This success can be attributed to the enormous capacity of multi-layer neural networks in modeling complex distributions of image and text data as well as the intricate design of the training mechanisms in GANs and diffusion models. The promising results of these frameworks have inspired the development of several methodologies for the training and evaluation of generative models in the literature.
@@
While the existing literature on deep generative models has mostly focused on centralized settings with training data stored by a single learner, many modern applications of deep learning algorithms are aimed at distributed scenarios where training data are collected by multiple agents in a network. A well-known instance of such a distributed setting is the federated learning task (McMahan et al., 2017), where several clients are connected to a server and aim to train a decentralized model through their communications with the server node while preserving the privacy of their collected data. A significant challenge in such distributed learning settings is the heterogeneous data distributions across clients, since the background features of every client could lead to a different data distribution. Especially, in training a deep generative model over a distributed network, the heterogeneity of the clients’ distributions could highly impact the performance and evaluation of the trained model.
@@
Figure 1: Centralized (score-all) vs. distributed (score-avg) evaluation of generative models. The centralized and distributed FID-based evaluations can lead to different rankings, while we prove the KID-based evaluations result in the same ranking.
In this work, we focus on the evaluation of deep generative models in heterogeneous distributed learning settings. Our primary goal is to highlight the challenges of extending standard evaluation metrics for generative models from the centralized setting to the heterogeneous distributed case. To do this, we consider and analyze the following two sensible extensions of an evaluation score: 1) the average score over clients (score-avg), i.e. the mean of the evaluation scores for individual clients, 2) the score with respect to the aggregate distribution of all clients (score-all) considering all the samples in the network. We consider standard assessment scores and analyze the consistency of model rankings suggested by the described extensions of the score to the distributed learning problem.
While the described score-all aggregation has been commonly used for the evaluation of generative models in the existing literature on distributed and federated generative model learning, we note that the estimation of the score-all metric will be challenging in a real distributed learning scenario. These challenges are mainly due to the privacy and computation constraints in standard distributed learning problems, preventing clients from sharing their observed samples and any revealing statistics of their training data with the rest of the network. On the other hand, the score-avg can be computed more efficiently since it only requires the client-based score value, which needs minor information on the samples and low communication expenses. However, it remains unclear if the original ranking of generative models according to score-all would be preserved under the score-avg assessment, because standard distances used for the evaluation task are non-linear in the input distributions.
In our theoretical analysis, we specifically focus on the Fréchet inception distance (FID) (Heusel et al., 2017) and kernel inception distance (KID) (Bińkowski et al., 2018) scores, which have been widely used in the evaluation of generative models. For the FID score, not only do we prove that the FID-all and FID-avg take different values, but further we show that the model rankings according to the two scores can be different. Indeed, we prove that the generative models with the optimal FID-avg and FID-all scores are different under heterogeneous client distributions, revealing the discrepancy between the two aggregate assessments.
On the other hand, in the case of KID score we prove that the aggregate scores KID-all and KID-avg result in the same rankings of generative models although they could take different values. This result shows the consistency of the KID evaluation between the mean aggregation in KID-avg and the collective data-based aggregation in KID-all. Therefore, our theoretical results suggest the different consistency behaviors of FID and KID scores when aggregated over a heterogeneous distributed learning setting: while averaging individual KID scores results in the same ranking as the KID score with respect to the distribution of collective client’s data, the same conclusion does not hold for FID and more generally Wasserstein distance-based evaluation metrics.
Finally, we discuss several numerical results on standard image datasets and generative model architectures which support our theoretical comparison of the aggregate evaluation scores in federated learning problems. Our empirical results suggest that the FID-all and FID-avg aggregations could lead to inconsistent rankings of the trained generative models in standard GANs and diffusion models. In our experiments, we observed that FID-avg often takes higher values when generated samples look sharper. On the other hand, the FID-all score seemed to be higher scores for models with higher diversity in their generated samples. In contrast, KID-all and KID-avg resulted in consistent rankings of the trained models where the evaluated score seems to aggregate both quality and diversity performance. In our numerical experiments, we also evaluated the precision, recall (Sajjadi et al., 2018; Kynkäänniemi et al., 2019), density, and coverage (Naeem et al., 2020) scores under the two introduced aggregations which, except in the case of recall score, could lead to different rankings of the trained model under the discussed score aggregations. In the following, we summarize the main contributions of our study:
  • Highlighting the challenges of evaluating generative models in heterogeneous distributed learning settings;
  • Analyzing two types of aggregate FID scores in distributed learning scenarios and proving the inconsistencies between the optimal models under the two scores;
  • Demonstrating the consistent rankings suggested by the client-wise averaged KID and the aggregate-data-based KID evaluations;
@@ 2 Related Work
A large body of related works (Theis et al., 2016) has focused on the evaluation of generative models in standard centralized learning settings. The existing evaluation scores can be categorized into two general groups: 1) Distance-based metrics defining a distance between the distribution of training data and learnt generative model. The distance between the real and fake distributions is usually computed after passing the samples through a pre-trained neural net offering a proper embedding of image data. The well-known evaluation scores in this category are the FID (Heusel et al., 2017) and KID (Bińkowski et al., 2018) scores. 2) Quality and diversity-based scores which output a score based on the sharpness and variety of the generated samples. The widely-used evaluation metrics in this category are the Inception score (Salimans et al., 2016), precision and recall metrics (Sajjadi et al., 2018; Kynkäänniemi et al., 2019), and density and coverage scores (Naeem et al., 2020). We note that in this work our goal is not to introduce a new evaluation metric, and our aim is to analyze the extensions of these scores to distributed and federated learning problems under heterogeneous data distributions across clients.
In another set of related works, extensions of generative model training methods including GANs and diffusion models to distributed federated learning have been studied. (Rasouli et al., 2020) propose Fed-GAN to train GANs in a federated learning setting. In (Hardy et al., 2019), the gradient from sample generation for the generator is exchanged on the server, while each client possesses a personalized discriminator. According to (Yonetani et al., 2019), different weights are assigned to local discriminators in the non-i.i.d. setting. Conversely, (Wu et al., 2022) ensure client privacy by sharing the discriminator across clients, while keeping the generator private. Additionally, (Su et al., 2023) explore a dual diffusion paradigm to extend diffusion-based models into the federated learning setting, addressing concerns related to data leakage.
@@ 3 Preliminaries
3.1 Deep Generative Models
In a deep generative model framework, a neural network generator $G$ is used to map a hidden random vector $\mathbf{Z}\in\mathbb{R}^{d}$ drawn according to a fixed distribution, e.g. isotropic multivariate Gaussian $\mathcal{N}(\mathbf{0},\sigma^{2}I)$ to a real-like sample $G(\mathbf{Z})$. Several deep learning approaches have been proposed to train such a generator network, including maximum-likelihood-based methods such as variational autoencoder (Kingma & Welling, 2013) and flow-based models (Dinh et al., 2016), generative adversarial networks (GANs) (Goodfellow et al., 2014), and denoising diffusion models (Ho et al., 2020). In our numerical evaluation, we mainly concentrated on the latter two methods, GANs and diffusion models, due to their state-of-the-art performance in computer vision applications.
In GANs, the training of generative models is framed as a min-max game between a generator network $G$ mapping latent vector $\mathbf{Z}$ to a real-like output and a discriminator network $D$ attempting to differentiate $G$’s generated samples from real training data. The GAN game is typically formulated as the following min-max optimization problem where $\bm{\theta},\,\bm{\omega}$ represent the parameters of generator and discriminator neural nets, and $f\bigl{(}G_{\bm{\theta}},D_{\bm{\omega}}\bigr{)}$ is the min-max objective representing $D$’s dissimilarity score for the generated and real samples:
$\min_{\bm{\theta}}\;\max_{\bm{\omega}}\;f\bigl{(}G_{\bm{\theta}},D_{\bm{\omega}}\bigr{)}.$ (1)
The training of GANs in a distributed learning problem aims at solving the above problem via a distributed optimization method. For example, in a federated learning setting where the local clients are connected to a single server node, the training of GAN players can be achieved by a federated min-max optimization algorithm as discussed in the related work section.
In the case of diffusion models, the generative model performs by multi-step denoising of a Gaussian input. The training of this approach is typically done by reversing the denoising process where the training data are turned to a Gaussian input via an iterative addition of independent Gaussian noise vectors. To extend diffusion models in federated learning, we follow the simple FedAvg (McMahan et al., 2017) method and average the locally updated diffusion networks at the server followed by synchronizing the clients with the averaged model.
3.2 Distance-based Evaluation of Generative Models
@@
$\mathrm{FID}(P_{X},P_{G})\>:=\>$ $\bigl{\|}\bm{\mu}_{X}-\bm{\mu}_{G}\bigr{\|}^{2}_{2}$
$\>+\mathrm{Tr}\Bigl{(}C_{X}+C_{G}-2\bigl{(}C_{X}C_{G}\bigr{)}^{1/2}\Bigr{)}.$
@@
$\;\mathrm{KID}\bigl{(}P_{X},P_{G}\bigr{)}:=\;$
$\sup_{\begin{subarray}{c}f\in\mathcal{H}_{k}:\,\|f\|_{\mathcal{H}_{k}}\leq 1\end{subarray}}\Bigl{(}\mathbb{E}_{X\sim P_{X}}\bigl{[}f(X)\bigr{]}-\mathbb{E}_{X^{\prime}\sim P_{G}}\bigl{[}f(X^{\prime})\bigr{]}\Bigr{)}^{2}$
@@
4 Evaluation of Generative Models in Distributed Learning Settings
In this section, we discuss two extensions of distance-based evaluation scores from a centralized case to heterogeneous distributed learning settings. In our analysis, we use $\mathcal{D}(P_{X},P_{G})$ to denote a general distance between data distribution $P_{X}$ and the generative model $P_{G}$. For example, $\mathcal{D}$ can be chosen to be the FID score or KID score, which we will analyze later in the section.
In a standard centralized setting, we have only a single distribution $P_{X}$ for real data. However, the main characteristic of a heterogeneous distributed learning problem is the multiplicity of the involved clients’ distribution. Here, we suppose a distributed setting $k$ clients and use $P_{X_{1}},\ldots,P_{X_{k}}$ to denote their underlying distributions, i.e. $P_{X_{i}}$ stands for the data distribution at client $i$. In addition, we assume that every client $i$ contributes a fraction $10<\lambda_{i}<$ 1 of the data in the network, that is $\lambda_{i}=\frac{n_{i}}{n}$ with $n=\sum_{j=1}^{k}n_{j}$ is the total number of samples in the network and $n_{i}$ is the number of samples at client $i$.
@@ 4 Evaluation of Generative Models in Distributed Learning Settings
  2. 2.
In the above, note that $\sum_{i=1}^{k}\lambda_{i}P_{X_{i}}$ is indeed a mixture distribution with $k$ components $P_{X_{1}},\ldots,P_{X_{k}}$ with frequency weights $\lambda_{1},\ldots,\lambda_{k}$. To relate the above aggregate scores, we first observe that when $\mathcal{D}$ is a convex function of the input distributions, which applies to both FID and KID scores, the score-avg $\mathcal{D}_{\mathrm{avg}}$ will upper-bound the score-all $\mathcal{D}_{\mathrm{all}}$:
Observation 1.
If $\mathcal{D}(P_{X},P_{G})$ is a convex function of $P_{X}$, then
$\mathcal{D}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\leq\mathcal{D}_{\mathrm{avg}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}$
Remark 1.
The convexity assumption on distance $D$ in the above observation applies to standard divergence scores, including Wasserstein distances, $f$-divergence measures, total variation distance, and the maximum mean discrepancy. Consequently, the result applies to the FID and KID scores.
While the mentioned observation shows how the two aggregate scores are compared with one another, it does not imply a monotonic relationship between $\mathcal{D}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}$ and $\mathcal{D}_{\mathrm{avg}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}$. Therefore, this observation does not provide a comparison of the ranking of generative models according to the two aggregate scores. In the following subsections, we study this question for the particular FID and KID scores.
4.1 Aggregate FID Scores in Distributed Learning
In the case of FID score, we utilize the formulation of FID as the 2-Wassesrtein-distance between the Gaussian-fitted model that leads to a Riemannian geometry. This observation results in the following theorem on FID-all and FID-avg aggregations. We defer the proofs to the Appendix.
@@ 4.1 Aggregate FID Scores in Distributed Learning
Suppose that $P_{X_{1}},\ldots,P_{X_{k}}$ are the clients’ distributions with the mean parameters $\bm{\mu}_{1},\ldots,\bm{\mu}_{k}$ and covariance matrices $C_{1},\ldots,C_{k}$, respectively, in the semantic space of the Inception net. Then, the following holds for a generative model $P_{G}$ with mean $\bm{\mu}_{G}$ and covariance $C_{G}$.

### v4

# On the Distributed Evaluation of Generative Models

@@ Abstract
The evaluation of deep generative models has been extensively studied in the centralized setting, where the reference data are drawn from a single probability distribution. On the other hand, several applications of generative models concern distributed settings, e.g. the federated learning setting, where the reference data for conducting evaluation are provided by several clients in a network. In this paper, we study the evaluation of generative models in such distributed contexts with potentially heterogeneous data distributions across clients. We focus on the widely-used distance-based evaluation metrics, Fréchet Inception Distance (FID) and Kernel Inception Distance (KID). In the case of KID metric, we prove that scoring a group of generative models using the clients’ averaged KID score will result in the same ranking as that of a centralized KID evaluation over a collective reference set containing all the clients’ data. In contrast, we show the same result does not apply to the FID-based evaluation. We provide examples in which two generative models are assigned the same FID score by each client in a distributed setting, while the centralized FID scores of the two models are significantly different. We perform several numerical experiments on standard image datasets and generative models to support our theoretical results on the distributed evaluation of generative models using FID and KID scores.
@@ 1 Introduction
Deep generative models including variational autoencoders [1], generative adversarial networks (GANs) [2], and diffusion models [3] have attained remarkable results over a wide array of machine learning tasks. These frameworks’ success can be attributed to the large capacity of multi-layer neural networks in modeling complex distributions of image and text data as well as the intricate design of the training mechanisms in these algorithms. The promising results of these generative modeling schemes have inspired the development of several assessment methods to properly compare and rank a group of generative models using their produced data.
A standard approach to evaluating generative models is to compare the distribution of their generated data with the distribution of samples in a reference set, typically chosen to be the test sample set in a learning problem. This comparison is often made using a distance measure between the generative and reference distributions. Two standard distance-based metrics are Fréchet Inception Distance (FID) [4] and Kernel Inception Distance (KID) [5]. In a standard centralized comparison of generative models, only a single reference distribution is considered for the evaluation. Therefore, to rank a group of generative models in a centralized setting, one can sort their distances to the reference distribution.
@@
On the other hand, several modern applications of deep generative models concern distributed settings where the reference data are provided by multiple clients in a network. A well-known distributed scenario is the federated learning setting [6], where several clients are connected to a server and aim to train and assess a decentralized model through their communications with the server node while preserving the privacy of their collected data. An important consideration for conducting evaluation in distributed settings is the potential heterogeneous data distributions across clients, since the background features of every client could lead to a different data distribution. For example, as displayed in Figure 1, the distribution of photos of clients in a mobile network could be highly diverse as they could possess different interests and characteristics.
@@
Figure 1: Centralized (score-all) vs. distributed (score-avg) evaluation of generative models. The centralized and distributed FID-based evaluations can lead to different rankings, while we prove the KID-based evaluations result in the same ranking of a group of generative models.
In this work, we focus on evaluating generative models in distributed settings and attempt to study the assessment task’s complexity under diverse client data distributions. Following the existing literature on training generative models in distributed settings [7, 8, 9, 10, 11, 12], the ultimate goal is to perform a centralized evaluation where the reference set is the collective dataset of all clients’ data. In our analysis, we call the evaluation score following such a centralized approach the score-all metric, which means the score is computed considering a reference set with all the clients’ data. However, the commonly considered privacy constraints in distributed settings restrict the use of such an evaluation method, because the assessment process requires the clients to share their data with the server.
Another sensible approach to private evaluation in distributed settings is to query every client about her own evaluation score for the generative model, and then to aggregate the clients’ individual evaluation scores. While this aggregation-based approach needs every client to share only her score, the resulting ranking of generative models could be different from the centralized score-all evaluation. Therefore, a key question is how consistent a score aggregation-based evaluation could be with the score-all assessment.
In this work, we answer the mentioned question for the KID and FID evaluation metrics. In the KID case, we prove that accessing to the clients’ individual KID scores will provide sufficient information for ranking a group of generative models using the centralized KID-all metric. Specifically, we prove that the mean of clients’ KID scores, which we call KID-avg, will result in the same ranking of generative models as the potentially expensive-to-compute KID-all score. This result reveals the low communication and privacy costs of ranking the models according to the centralized KID-all score, which is identical to the ranking of the generative models using the averaged KID-avg score.
In contrast, we show that the knowledge of clients’ individual FID scores for a set of generative models is insufficient to compare those models according to the centralized FID-all score. We provide general examples of distributed scenarios in which the clients report the same FID values for two generative models, while the models lead to significantly different centralized FID scores. Furthermore, we show that the averaged FID scores across users, which we call the FID-avg score, would be minimized by a different generative model from the centralized-all minimizing model. Therefore, our theoretical results indicate the insufficiency of the clients’ individual FID scores for ranking the models according to their centralized FID-all score.
Finally, we discuss several numerical results on standard image datasets and generative model architectures which are consistent with our analytical results on the aggregation of the distance-based evaluation scores. We utilize the truncation technique to simulate a heterogeneous federated learning scenario where the clients’ image data have different distributions, and show several cases where two models share the same client-based FID scores, while their FID with respect to the collective set of clients’ data is significantly different. We also show the inconsistency between the FID-all and FID-avg’s rankings of a group of generative models. On the other hand, Our empirical results suggest that the centralized KID-all and distributed FID-avg aggregations provide consistent rankings of the tested generative models, while they could have different values. In the following, we summarize the main contributions of our study:
  • Studying the distance-based evaluation of generative models in heterogeneous distributed settings,
  • Proving the identical rankings given by the centralized KID-all score with the clients’ collective reference set and the KID-avg score averaging clients’ individual KIDs.
  • Demonstrating the insufficient knowledge of clients’ FID scores for ranking generative models based on the FID-all score with the clients’ collective reference dataset;
@@ 2 Related Work
A large body of related works [13] has focused on the evaluation of generative models in standard centralized learning settings. The existing evaluation scores can be categorized into two general groups: 1) Distance-based metrics defining a distance between the distribution of training data and learnt generative model. The distance between the real and fake distributions is usually computed after passing the samples through a pre-trained neural net offering a proper embedding of image data. The well-known evaluation scores in this category are the FID [4] and KID [5] scores. 2) Quality and diversity-based scores which output a score based on the sharpness and variety of the generated samples. The widely-used evaluation metrics in this category are the Inception score [14], precision and recall metrics [15, 16], and density and coverage scores [17]. We note that in this work our goal is not to introduce a new evaluation metric, and our aim is to analyze the extensions of these scores to distributed and federated learning problems under heterogeneous data distributions across clients.
In another set of related works, extensions of generative model training methods including GANs and diffusion models to distributed federated learning have been studied. [7] propose Fed-GAN to train GANs in a federated learning setting. In [9], the gradient from sample generation for the generator is exchanged on the server, while each client possesses a personalized discriminator. According to [10], different weights are assigned to local discriminators in the non-i.i.d. setting. Conversely, [11] ensure client privacy by sharing the discriminator across clients, while keeping the generator private. Additionally, [12] explore a dual diffusion paradigm to extend diffusion-based models into the federated learning setting, addressing concerns related to data leakage.
@@ 3 Preliminaries
$\mathrm{FID}(P_{X},P_{G})\>:=\>$ $\bigl{\|}\bm{\mu}_{X}-\bm{\mu}_{G}\bigr{\|}^{2}_{2}\>+\mathrm{Tr}\Bigl{(}C_{X}+C_{G}-2\bigl{(}C_{X}C_{G}\bigr{)}^{1/2}\Bigr{)}.$
@@
$\;\mathrm{KID}\bigl{(}P_{X},P_{G}\bigr{)}:=\;\sup_{\begin{subarray}{c}f\in\mathcal{H}_{k}:\,\|f\|_{\mathcal{H}_{k}}\leq 1\end{subarray}}\Bigl{(}\mathbb{E}_{X\sim P_{X}}\bigl{[}f(X)\bigr{]}-\mathbb{E}_{X^{\prime}\sim P_{G}}\bigl{[}f(X^{\prime})\bigr{]}\Bigr{)}^{2}$
@@
4 Evaluation of Generative Models in Distributed Contexts
In this section, we discuss two extensions of distance-based evaluation scores from a centralized case to heterogeneous distributed learning settings. In our analysis, we use $\mathcal{D}(P_{X},P_{G})$ to denote a general distance between data distribution $P_{X}$ and the generative model $P_{G}$. Considering the standard distance-based metrics, $\mathcal{D}$ can be chosen to be the FID or KID score, which we will analyze in this section.
In a standard centralized setting, we have only a single distribution $P_{X}$ for real data. However, the main characteristic of a heterogeneous distributed learning problem is the multiplicity of the involved clients’ distribution. Here, we suppose a distributed setting with $k$ clients and use $P_{X_{1}},\ldots,P_{X_{k}}$ to denote their underlying distributions, i.e. $P_{X_{i}}$ stands for the data distribution at client $i$. In addition, we assume that every client $i$ has a fraction $0\leq\lambda_{i}\leq 1$ of the data in the network, i.e, $\lambda_{i}=\frac{n_{i}}{n}$ where $n=\sum_{j=1}^{k}n_{j}$ is the total number of samples in the network and $n_{i}$ is the number of samples at client $i$.
@@ 4 Evaluation of Generative Models in Distributed Contexts
In the above, note that $\sum_{i=1}^{k}\lambda_{i}P_{X_{i}}$ is indeed a mixture distribution with $k$ components $P_{X_{1}},\ldots,P_{X_{k}}$ with frequency parameters $\lambda_{1},\ldots,\lambda_{k}$. As discussed in the introduction, the score-all metric is the desired evaluation score following the literature on distributed generative modeling. However, due to the privacy and communication costs, the score-avg metric is cheaper to compute over the network. Therefore, understanding the connections between the rankings of the models following the two aggregate scores is the key to apply the score-avg as a surrogate for the score-all metric in the assessment process.
4.1 KID-based Evaluation in Distributed Settings
Here, we focus on the KID score and analyze the consistency of model rankings following KID-all and KID-avg scores. The following theorem proves that the KID-all and KID-avg are guaranteed to provide a consistent ordering of a group of generative models, and there is a monotonic relationship between the two scores independent of the choice of the generative model. We defer the theoretical statements’ proof to the Appendix.
@@ 4.1 KID-based Evaluation in Distributed Settings
$\mathrm{KID}_{\mathrm{avg}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\,=\,\mathrm{KID}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}+\sum_{i=1}^{k}\lambda_{i}\,\mathrm{KID}\bigl{(}\widehat{P}_{X},P_{X_{i}}\bigr{)},$
The above theorem shows the consistent rankings implied by the aggregate KID-avg and KID-all scores, as the difference between the scores remains constant while changing the model $P_{G}$. Therefore, to rank a set of models based on the desired KID-all score, the clients can only share their own evaluated KID, and the server can average the reported scores to recover the models’ ranking according to the centralized KID-all.
4.2 FID-based Evaluation in Distributed Settings
In the case of FID score, we utilize the formulation of FID as the 2-Wassesrtein-distance between the Gaussian-fitted model that leads to a Riemannian geometry. This observation results in the following theorem on FID-all and FID-avg aggregations.
Theorem 2.
Suppose that $P_{X_{1}},\ldots,P_{X_{k}}$ are the clients’ distributions with the mean parameters $\bm{\mu}_{1},\ldots,\bm{\mu}_{k}$ and covariance matrices $C_{1},\ldots,C_{k}$, respectively, in the Inception net embedding. Then, the following holds for a generative model $P_{G}$ with mean $\bm{\mu}_{G}$ and covariance $C_{G}$.